A certain 6% annual coupon rate convertible bond (maturing in 20 years) is convertible at the holders option into 20 shares of common stock. The bond is currently trading at $800. the stock (which pays .71 a share in annual dividens) is currently priced in tje market at $29.78 a share. a) what is the bonds conversion price b) what is the conversion ratio c) what is the conversion value & parity d) what is the conversion permium in dollars and as a percentage e) what is the bonds payback period f) if comparably rated, nonconvertible bonds sell to a yeild of 8%, what is the investment value of the convertible?
Answer #1

Ans - (a) The Conversion Price of the bond will be calculated by the formula

Bond Conversion Price = Price of Bond/No of common shares=$800/20= $40.

Ans-(b) Bonds conversion ratio = Price of Bonds/ Bonds conversion Price = $800/40= 20. Since we got the Bonds conversion Price as $40.

Ans- (c) In part c we have to calculate the conversion value and conversion parity.

Conversion value is given by the multiplication of conversion ratio and a current stock price of the bond.

Therefore the conversion value = Current stock price of bond*converstion ratio = $29.78*20= $595.6 is the conversion value.

Now the conversion parity is given by the price of the convertible bond price divided by the number of common shares. Therefore the conversion parity will be $800/20 = $40.

Ans- (d) Conversion Premium is given by the difference of current bond price and the conversion value. Therefore the conversion premium will be Bond price - conversion value = $800 - $595.6 = $204.4.is the conversion premium in dollars.

Conversion Premium Percentage is given by conversion premium/conversion value*100 =204/595.6*100=34.25%.

Note:- As per Chegg rule, we are allowed to give 4 answers.